OpenStudy (anonymous):

\int _{ 0 }^{ pi }{ \frac { sin\theta }{ \sqrt { pi\quad -\quad \theta } } } using\quad \frac { 1 }{ \sqrt { pi\quad -\quad \theta } } \quad as\quad comparison

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

\[\int\limits _{ 0 }^{ \pi }{ \frac { \sin\theta }{ \sqrt { \pi\quad -\quad \theta } } } using\quad \frac { 1 }{ \sqrt { \pi\quad -\quad \theta } } \quad as\quad comparison\]

OpenStudy (anonymous):

sorry idk how to do it, good luck though!

OpenStudy (anonymous):

Hahaha, thanks.

OpenStudy (sirm3d):

you're asking if the integral is convergent or divergent?

OpenStudy (anonymous):

Yeah

OpenStudy (sirm3d):

if \[\int_0^{\pi}\frac{d\theta}{\sqrt{\pi-\theta}}\] is convergent, you can conclude that the integral in question is also convergent.

OpenStudy (anonymous):

How do I know that integral you stated above is convergent? That's the part that is troubling me.

OpenStudy (sirm3d):

why don't you evaluate the definite integral? if it is a value, then the integral is convergent.

OpenStudy (anonymous):

Oh okay, I thought it was improper. I just realized that it was definite. Thanks!

OpenStudy (sirm3d):

it is an improper integral, but after integration, the function is continuous in that interval, so you can treat it like an ordinary definite integral.